OpenStudy (anonymous):

Julie and eric row their boat ( at a constant speed) 40 miles downstream for 4 hours,helped by the current. Rowing at the same rate the trip back against the current takes 10 hours, find the rate of the current \

OpenStudy (kropot72):

Let x = current rate On down stream trip rowers rate = 40/4 - x On th eupstream trip the overall rate 40/10 = (40/4) - x -x Solving for x gives the rate of the current as 3 mph

OpenStudy (anonymous):

this stuff is hard for me, its like a foreign language. Thanks for your help. can you break it down for me though? How did you get 3 I see the equation and still dont get it

OpenStudy (anonymous):

sorry

OpenStudy (kropot72):

np. The overall rate on the upstream trip is the rowers rate (given by 40/4 mph) minus the opposing current rate (given as x):\[\frac{40}{10}=(\frac{40}{4}-x)-x\] This simplifies to:\[4=10-x-x=10-2x\] Add 2x to both sides and subtract 4 from both sides to get: \[2x=10-4=6\] \[x=\frac{6}{2}=3\]

OpenStudy (anonymous):

okay thats better , thanks your a life saver!!!

OpenStudy (kropot72):

Gr8. You're welcome :)